# thanks: Corresponding author

### v1

@@ Abstract
Förster resonance energy transfer (FRET) is a quantum mechanical phenomenon involving the non-radiative transfer of energy between coupled electric dipoles. Due to the strong dependence of FRET on the distance between the dipoles, it is frequently used as a “molecular ruler” in biology, chemistry, and physics. This is done by placing dipolar molecules called dyes on molecules of interest. In time-resolved confocal single-molecule FRET (smFRET) experiments, the joint distribution of the FRET efficiency and the donor fluorescence lifetime can reveal underlying molecular conformational dynamics via deviation from their theoretical Förster relationship. This deviation is referred to as a dynamic shift. Quantifying the dynamic shift caused by the motion of the fluorescent dyes is essential to decoupling the dynamics of the studied molecules and the dyes. We develop novel Langevin models for the dye linker dynamics, including rotational dynamics, based on first physics principles and proper dye linker chemistry to match accessible volumes predicted by molecular dynamics simulations. By simulating the dyes’ stochastic translational and rotational dynamics, we show that the observed dynamic shift can largely be attributed to the mutual orientational dynamics of the electric dipole moments associated with the dyes, not their accessible volume.
FRET, Langevin Dynamics, Kappa Squared, Time-inhomogeneity, Dynamic shift
@@ I Introduction
Förster resonance energy transfer (FRET)[1, 2] experiments are commonly used in biochemistry and biophysics to measure distances on molecules of interest.[3, 4, 5, 6] This is done by exploiting the physical mechanism of FRET wherein energy is transferred non-radiatively between a donor and an acceptor fluorophore due to electric dipole coupling. Throughout the document, these fluorophores will be referred to as dyes. The rate of energy transfer depends on the distance between the dipoles to the sixth power, as well as an orientational factor and several other time-independent factors.[1, 4, 7, 8, 9] Due to this strong dependence on distance, FRET is often used as a molecular ruler [10, 11]. Therefore, an accurate understanding of FRET measurements is paramount to understanding the collected experimental data and forming concrete hypotheses about the conformational dynamics of the molecule of interest, which provides information for the health sciences. [12]
It has been shown in [13, 14, 15] that the joint distribution between FRET efficiency and lifetime measurements contains a wealth of information on the structural dynamics of the underlying molecule. This information is seen by a deviation from the ideal Förster relationship, termed a dynamic shift [13]. The use of smFRET in structural biology is well established. [16, 17, 18, 19, 11, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 14, 31, 32, 33, 19, 18] Moreover, the use of the dynamic shift to understand dynamic structural biology has recently been established. [30, 14, 31, 34, 35, 21] Due to thermal noise, the dye position will fluctuate randomly, inducing a change in the joint FRET-lifetime distribution.[36, 15] Hence, it is necessary to understand the influence of this stochastic fluctuation on the resulting FRET measurements.[37, 38, 39, 40, 41] So, a complete understanding of the dynamic shift induced by the dye dynamics is needed to facilitate accurate estimates for structural dynamics.[42, 43]
Currently, in the literature, there is no physical model for the dye motion to facilitate accurate and high-precision FRET measurements.[36, 15] The models available are either simplistic [44] or require a high computational cost [20, 45, 43, 46]. Typically, all-atom molecular dynamics simulations are conducted on a sample-by-sample basis to calculate the accessible volume of the dyes in order to estimate the range of motion of the molecule and, therefore, provide a way of uncertainty quantification in the FRET-lifetime distribution [43, 34, 31, 47]. An important critique of this method is that the simulations do not last for the entire sampling process.[43] Hence, the full range of motion for the dyes during a sample time is not achieved. Here, we attempt to provide a semi-analytic model for fluorescent dye motion. In addition, these simulations are meant to approximate the equilibrium distribution of the dye motion and hence ignore any time inhomogeneity in the energy transfer process itself. We consider several of the following relevant aspects of dye motion. First, will an isotropic Gaussian process provide a correct model? In this situation, the common assumption of Gaussian inter-dye displacement is satisfied. Second, what is the impact of linker length? Third, what is the role of dipole motion in FRET measurements? We answer these questions in the context of a simulated smFRET experiment. Our results shed light on the past assumption that the dynamic shift induced by the dyes is due to the accessible volume of the dye. [13, 14] In reality, the dynamic shift has to do with the full state dynamics of the dyes, both translational and orientational.
@@ I.1 Confocal smFRET
I.2 FRET Model as a Continuous-Time Markov Chain (CTMC)
Consider two completely static dyes with normalized dipole moments $\mu_{A}\in S^{2}$ and $\mu_{D}\in S^{2}$ for the acceptor and donor, respectively. Further, let the inter-dye displacement vector be $r\in\mathbb{R}^{3}$. The energy transfer rate is defined in Eq. (1),
@@ I.2 FRET Model as a Continuous-Time Markov Chain (CTMC)
where $k_{D}$ is the fluorescence decay time for the donor dye, and $R_{0}$ is a physical parameter called the Förster radius where the probability of energy transfer is $0.5$ [2, 17, 4, 41]. Note that the energy transfer rate increases dramatically as the distance decreases and vice versa as the distance increases. However, no matter the distance, no energy transfer can happen if $\mu_{D},\mu_{A}$, and $\hat{r}$ are mutually orthogonal. [49, 50] The Förster radius can be written as $R_{0}^{6}(t)=C\kappa^{2}(t)$ where $C$ is a constant depending on the environment surrounding the dye. The parameter $\kappa^{2}(t)$ is the dipole orientational factor
$\kappa(t)=(\mu_{D}(t)\cdot\mu_{A}(t))-3(\hat{r}\cdot\mu_{D}(t))(\hat{r}\cdot\mu_{A}(t))$ (2)
@@ I.3 The Dynamic Shift
Figure 2: Visualization of the definition of the dynamic shift using normalized values: $S=\{(\mathcal{E},\tau)\in(0,1)^{2}:\mathcal{E}+\tau-1=0\}$ in the bottom. The top figure is a visualization of the moment difference dynamic shift. For the moment difference We used $\mathcal{E}_{\tau}=1-\frac{\overline{\tau_{D(A)}}}{\tau_{D(0)}}$
@@
The dynamic shift introduced in [13] considers an underlying distribution dependent on two separate states. Consider two FRET efficiency states denoted by $\mathcal{E}_{i},~{}i=1,2$ with equal transition rate between the states $\lambda$ for simplicity. Such a two-state system provides valuable insight into the nature of the dynamic shift. When two states are separated on long time scales, $\lambda<<1$, the dynamic shift is slight due to the small amount of mixing during a burst or sample. As the two states mix, corresponding to an increase in $\lambda$, an arc forms between the static FRET-lifetime coordinates, following $(1-\mathcal{E}_{1}-\mathcal{E}_{2})\mathcal{E}-\mathcal{E}_{1}\mathcal{E}_{2}$. As $\lambda\rightarrow\infty$, this process culminates in a point mass FRET-lifetime distribution with a dynamic shift at the maximum of this arc. Therefore, the dynamic shift can be seen as a metric of the amount of mixing between states. Two-state transition systems can be used to understand the transition rates between stable states in protein conformational dynamics. For the current purpose, it provides a convenient method for interpreting the dynamic shift induced by the dyes. The dynamic shift will most readily be present when there are mixing states, and understanding the influence of mixing between an uncountable number of states is of current interest. It will be shown in Section III that the dynamic shift induced by dye dynamics can be viewed as a consequence of the fluctuations in the energy transfer rate during the FRET process. Under common circumstances, the energy transfer rate can be approximated by a two-state system corresponding to the modes of the distribution, essentially leading to a quickly transitioning two-state system.
@@ II Stochastic Models of Fluorescence Dynamics
Throughout, the dynamics are assumed to evolve on different timescales. Letting $T_{P}$, $T_{D}$, $T_{O}$ represent the timescales of protein dynamics, dye translational dynamics, and dipole orientational dynamics. The order of timescale separation assumed in this work is given by $T_{P}>>T_{D}>>T_{O}$. Further, as in [53], it is assumed that the orientation process and the translational process are independent processes. Note that this is an extremely common assumption since the independence of $\kappa$ and $r$ dynamics is implicitly assumed whenever the average $\kappa$ value is used and whenever static $\kappa$ distributions are employed. [68] Moreover, to provide a clear and succinct picture of the influence of dye dynamics on FRET measurements, the timescale $T_{P}$ is not considered in the current discussion. However, an extension of this analysis to include this timescale is in development.
@@ II.1 Spring Models
$d\mathbf{X}_{t}=K(\mathbf{X}_{t}-\mathbf{X}_{eq})dt+\sigma I\circ dB_{t},$ (12)
where $K=[k_{i,j}]_{i,j=1}^{3}$ is a matrix of spring constants. The notation $\circ dB_{t}$ denotes the use of Stratonovich integration.[71] We refer to systems such that the spring matrices can be written in the form $K=kI_{3\times 3}$, as isotropic springs. Otherwise, the system is called anisotropic.
@@
In the isotropic case, illustrated in Figure 4A, the spring matrix is given by $k_{eff}I_{3\times 3}$. This provides a symmetric three-dimensional Gaussian as the stationary distribution for the isotropic spring.[69] It can be seen in [70, 73] that the variance of this distribution will be given by $\Sigma=\frac{\sigma}{k_{eff}}I_{3\times 3}.$
Figure 4: Dye model trajectories and resulting FRET efficiency vs normalized mean fluorescence lifetime and moments difference for A-B) Isotropic Spring, C-D) Anisotropic Spring, E-F) Elastic Pendulum. G) Sample trajectory of a spherical Brownian motion on the unit sphere. Such processes are used to model the diffusion of the electric dipole moment. H) FRET efficiency vs normalized mean fluorescence lifetime and moments difference when the dipole orientation for donor and acceptor is included in the FRET process during the Elastic Pendulum model.
In the anisotropic case, illustrated in Figure 4C, we use a diagonal spring matrix with two entries being $pk_{eff}$ and the third being $k_{eff}$ with $p\in(0,1)$. Therefore, the stationary distribution is an ellipse with major axes determined by the entries of the spring matrix.
This planar freedom in the anisotropic case can be used to investigate the influence of the orientation of the stationary distribution on the resulting dynamic shift. Such a scenario is exemplified in the case when the planes formed by the major axes of each stationary ellipse are mutually orthogonal. Since the stationary distribution for the isotropic case is a sphere and is perfectly symmetric, this can only arise in the anisotropic case.
Furthermore, these models have the added benefit of having an analytical expression for the inter-dye displacement, espeically in the isotriopic case. Sine the coordinates will be Gaussian distributed the distance between them is simply Reighley distributed. [64] This distribution is unimodal, and therefore the only mixing present is due to the variance of the stationary distributions. This mixing is, therefore, strongly dependent on the flexibility of the dyes, and as seen above, the variances in their positions when modeling the spring constant according to dye linker chemistry are quite small.
@@ II.2 Elastic Pendulum Model
Importantly, the system is considered in spherical coordinates. The radial dynamics, $r_{t}$ evolve according to the spring dynamics explained in Section II.1, with slight alterations due to the change of coordinates. The angular parts of the motion are given by the standard nonlinear pendulum force in the polar direction $\theta_{t}$ and free diffusion in the azimuthal direction $\phi_{t}$. Figure 4 shows a sample dye trajectory.
@@
Moreover, this model presents a possible explanation for the dynamic shift induced by dye motion due to the non-Gaussian inter-dye distributions, as shown in
@@ II.3 Orientational Dynamics
The final consideration involves the orientational dynamics of the electric dipole moments of the dyes. As discussed in Section I, the Förster radius is dependent on the $\kappa$ parameter, which is dependent on the mutual orientations of the electric dipole moments $\mu_{A},\mu_{D}$ and the inter-dye displacement unit vector $\hat{R}$. Typically, $\kappa^{2}$ is taken as the mean value of $2/3$ when the unit vectors are considered uniformly distributed on the sphere.[50, 68, 49] This assumption ignores the temporal aspect of the fluorescent process. Since dyes reorient on timescales faster than fluorescent lifetimes, the energy exchange rate changes during the FRET process. This changes the original CTMC model to a time-inhomogeneous CTMC, and thus, the transfer rates are dependent on the time integral of the infinitesimal transfer rates. [56, 57, 58, 55]
@@
$\begin{cases}\mathbf{I}d\omega_{t}+\omega_{t}\times\mathbf{I}\omega_{t}=-\nu\omega_{t}+dW_{t}\\ \omega_{t}=d\Phi_{t}\end{cases}$ (14)
where $dW_{t}$ is a spherical Brownian motion and $\Phi_{t}$ is the angular position vector. Assuming the dye is overdamped and hence $d\omega_{t}=0$, one obtains the simplified equations
$\begin{cases}\omega_{t}\times\mathbf{I}\omega_{t}=-\nu\omega_{t}+dW_{t}\\ \omega_{t}=d\Phi_{t}.\end{cases}$ (15)
@@
$\begin{cases}d\theta_{t}&=\frac{\sigma_{\theta}^{2}}{\tan{(\theta_{t})}}dt+\sigma_{\theta}\circ dB_{t}\\ d\phi_{t}&=\frac{\sigma_{\phi}}{\sin{\theta_{t}}}\circ dB_{t}\end{cases}$ (17)
The rotational diffusion coefficients depend on the hydrodynamic radius of the dye $R_{h}$ by the classical relation $D=8\pi\nu R_{h}^{3}$, where $\nu$ is the dynamic viscosity of the surrounding fluid. A sample trajectory is shown in Figure 3.
Note that the stationary distribution for such a system is the uniform distribution, providing an ideal starting stochastic process to test the time-dependent behavior of orientational dynamics.[71, 75] The key idea is that the excitation of the fluorophores provides a single initial $\kappa$ value. This value then needs to evolve back to equilibrium during the FRET process. The relaxation effects are the object of interest, especially with regard to lifetime duration. The notion that $\kappa$ may be close to 00 for the entire FRET process for one excitation but higher for another in the same sampling time provides an additional source of variance in the lifetime distribution. This should depend on the rotational diffusion of the dipole moment, with faster reorientation causing an averaging out effect as mentioned in [50].

### v2

@@ Abstract
Förster resonance energy transfer (FRET) is a quantum mechanical phenomenon involving the non-radiative transfer of energy between coupled electric dipoles. Due to the strong dependence of FRET on the distance between the dipoles, it is frequently used as a “molecular ruler” in biology, chemistry, and physics. This is done by placing dipolar molecules called dyes on molecules of interest. In time-resolved confocal single-molecule FRET (smFRET) experiments, the joint distribution of the FRET efficiency and the donor fluorescence lifetime can reveal underlying molecular conformational dynamics via deviation from their theoretical Förster relationship. This deviation is referred to as a dynamic shift. Quantifying the dynamic shift caused by the motion of the fluorescent dyes is essential to decoupling the dynamics of the studied molecules and the dyes. We develop novel Langevin models for the dye linker dynamics, including rotational dynamics, based on first principle physics and proper dye linker chemistry to match accessible volumes predicted by molecular dynamics simulations. By simulating the dyes’ stochastic translational and rotational dynamics, we show that the observed dynamic shift can largely be attributed to the mutual orientational dynamics of the electric dipole moments associated with the dyes, not their accessible volume. Our models provide the most up-to-date and accurate estimation of FRET.
FRET, Langevin Dynamics, Kappa Squared, Time-heterogeneity, Dynamic shift
@@ I Introduction
Förster resonance energy transfer (FRET)[1, 2] experiments are commonly used in biochemistry and biophysics to measure distances at the molecular level to resolve conformational states.[3, 4, 5, 6, 7] This is done by exploiting the physical mechanism of FRET wherein energy is transferred non-radiatively between a donor and an acceptor fluorophore due to electric dipole coupling. For simplicity, we refer to these fluorophores as dyes. In FRET, the rate of energy transfer depends on the distance between the dipoles to the sixth power, as well as an orientational factor and several other time-independent factors.[1, 5, 8, 9, 10] Due to this strong dependence on distance, FRET is often used as a molecular ruler [3, 11]. However, due to the many distance-related interpretations in FRET and potential pitfalls, assignments of FRET derived distances to specific structural properties is not trivial [12, 13, 14, 15, 8]. For example, measuring distances within single molecules (smFRET) is a well-established concept in structural biology.[16, 17, 18, 19, 11, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 19, 18] Hence, accurately understanding FRET measurements is crucial for interpreting experimental data and forming concrete hypotheses about the conformational dynamics of the molecule of interest. This understanding is essential for developing accurate biosensors, elucidating signal transduction pathways, studying biomolecular interactions, advancing drug development, and creating fluorescence spectroscopic toolkits. [35, 36]
FRET can be measured via photon counting or time-resolved methods.[37] Previously, we and others demonstrated [38, 31, 39] that the joint distribution between FRET efficiency—the proportion of excited donor molecules transferring energy to acceptor molecules—and fluorescence lifetime—the time a molecule remains in the excited state—provides extensive information on the structural dynamics of the underlying molecule. Thus, we introduced the term ”dynamic shift” to account for deviations from the ideal Förster relationship[38], with multiple exampled on how to the dynamic shift concept helped understand the structural dynamics of biomolecules.[30, 31, 32, 40, 41, 21] Due to thermal noise, the dye position fluctuates randomly, altering the joint FRET-lifetime distribution.[42, 39] Therefore, it is crucial to understand how these stochastic fluctuations influence the resulting FRET measurements.[43, 44, 45, 46, 47] A comprehensive understanding of the dynamic shift induced by dye dynamics is needed to achieve accurate estimates of structural dynamics.[48, 49]
Current dye models in the literature lead to significant uncertainties in FRET measurements.[42, 39] Moreover, the available models are either simplistic[50] or require high computational costs.[20, 51, 49, 52] For example, all-atom molecular dynamics (MD) simulations are often conducted on a sample-by-sample basis to calculate the accessible volume of the dyes, thereby estimating the range of motion of the dye and the labeled molecule while providing a method for uncertainty quantification in the FRET-lifetime distribution.[49, 53, 32, 54] However, a key critique of this method is that MD simulations do not last for the entire sampling process of typical smFRET experiments.[49] Additionally, MD simulations are designed to approximate the equilibrium distribution of the dye motion, ignoring any time heterogeneity in the energy transfer process itself. Consequently, the full range of dye motion during a sample period is not captured.
Here, we present a semi-analytic model for fluorescent dye motion that addresses the following questions: First, will an isotropic Gaussian process provide a correct model? Second, what is the impact of linker length? Third, what is the role of dipole motion in FRET measurements? We answer these questions in the context of simulated smFRET experiments. Our results challenge the previous assumption that the dynamic shift induced by the dyes is solely due to the accessible volume of the dye (i.e., translational motion).[38, 31] This work demonstrates that the dynamic shift actually depends on the full state dynamics of the dyes, including translational and orientational movements. Therefore, to reduce uncertainty in FRET-derived distances or determine the structural dynamics of biomolecules, we can decouple the individual dye dynamics from the dynamics of the molecule under study by their dynamic shift signature.
@@ I.1 Confocal smFRET
I.2 FRET Model
Consider two completely static dyes with normalized dipole moments $\hat{\mu}_{A}\in S^{2}$ and $\hat{\mu}_{D}\in S^{2}$ for the acceptor and donor, respectively. Further, let the inter-dye displacement vector be $r\in\mathbb{R}^{3}$. The energy transfer rate is defined in Eq. (1),
@@ I.2 FRET Model
where $k_{D}$ is the fluorescence decay rate for the donor dye, and $R_{0}$ is the distance at which the energy transfer efficiency is 0.5 [2, 17, 5, 47]. Note that the energy transfer rate increases steeply as the distance decreases and inversely as the distance increases. However, no matter the distance, no energy transfer can happen if $\hat{\mu}_{D},\hat{\mu}_{A}$, and $\hat{r}$ are mutually orthogonal. [56, 57] The Förster radius can be written as $R_{0}^{6}(t)=C\kappa^{2}(t)$ where $C$ is a constant depending on the environment surrounding the dye. The parameter $\kappa^{2}(t)$ is the dipole orientational factor
$\kappa(t)=(\hat{\mu}_{D}(t)\cdot\hat{\mu}_{A}(t))-3(\hat{r}\cdot\hat{\mu}_{D}(t))(\hat{r}\cdot\hat{\mu}_{A}(t))$ (2)
@@ I.3 The Dynamic Shift
Figure 2: Visualization of the definition of the dynamic shift using normalized values: $S=\{(\mathcal{E},\tau)\in(0,1)^{2}:\mathcal{E}+\tau-1=0\}$ in the bottom. The top figure is a visualization of the moment difference dynamic shift. For the moment difference we used $\mathcal{E}_{\tau}=1-\frac{\overline{\tau_{D(A)}}}{\tau_{D(0)}}$.
@@
The dynamic shift introduced in [38] considers an underlying distribution dependent on two separate states. Consider two FRET efficiency states denoted by $\mathcal{E}_{i},~{}i=1,2$ with equal transition rate between the states $\lambda$ for simplicity. Such a two-state system provides valuable insight into the nature of the dynamic shift. When two states are separated on long time scales, $\lambda<<1$, the dynamic shift is slight due to the small amount of mixing during a burst or sample. As the two states mix, corresponding to an increase in $\lambda$, an arc forms between the static FRET-lifetime coordinates, following $(1-\mathcal{E}_{1}-\mathcal{E}_{2})\mathcal{E}-\mathcal{E}_{1}\mathcal{E}_{2}$. As $\lambda\rightarrow\infty$, this process culminates in a point mass FRET-lifetime distribution with a dynamic shift at the maximum of this arc. Therefore, the dynamic shift can be seen as a metric of the amount of mixing between states. Two-state transition systems can be used to understand the transition rates between stable states in biomolecules conformational dynamics. For the current purpose, it provides a convenient method for interpreting the dynamic shift induced by the dyes. The dynamic shift will most readily be present when there are mixing states, and understanding the influence of mixing between an uncountable number of states is of current interest. It will be shown in Section III that the dynamic shift induced by dye dynamics can be viewed as a consequence of the fluctuations in the energy transfer rate during the FRET process. Under common circumstances, the energy transfer rate can be approximated by a two-state system corresponding to the modes of the distribution, essentially leading to a quickly transitioning two-state system.
@@ II Stochastic Models of Fluorescence Dynamics
Throughout, the dynamics are assumed to evolve on different timescales. Letting $T_{P}$, $T_{D}$, $T_{O}$ represent the timescales of biomolecules dynamics, dye translational dynamics, and dipole orientational dynamics. The order of timescale separation assumed in this work is given by $T_{P}>>T_{D}>>T_{O}$. Further, as in [60, 74, 75, 76, 77], it is assumed that the orientation process and the translational process are independent processes. Note that this is an extremely common assumption since the independence of $\kappa$ and $r$ dynamics is implicitly assumed whenever the average $\kappa$ value is used and whenever static $\kappa$ distributions are employed. [78] Moreover, to provide a clear and succinct picture of the influence of dye dynamics on FRET measurements, the timescale $T_{P}$ is not considered in the current discussion. However, an extension of this analysis to include this timescale is in development.
@@ II.1 Spring Models
$d\mathbf{X}_{t}=K(\mathbf{X}_{t}-\mathbf{X}_{eq})dt+\sigma\mathbb{I}\circ dB_{t},$ (12)
where $K=[k_{i,j}]_{i,j=1}^{3}$ is a matrix of spring constants and $\mathbb{I}$ is the identity matrix. The notation $\circ dB_{t}$ denotes the use of Stratonovich integration [81], where $B_{t}$ is Brownian motion. $\sigma>0$ is the volatility of the random fluctuations that are modeled as Brownian motions. We refer to systems such that the spring matrices can be written in the form $K=k\mathbb{I}_{3\times 3}$, as isotropic springs. Otherwise, the system is called anisotropic.
@@
In the isotropic case, illustrated in Figure 4A, the spring matrix is given by $k_{eff}\mathbb{I}_{3\times 3}$. This provides a symmetric three-dimensional Gaussian as the stationary distribution for the isotropic spring.[79] It can be seen in [80, 83] that the variance of this distribution will be given by $\Sigma=\frac{\sigma}{k_{eff}}\mathbb{I}_{3\times 3}.$
Figure 4: Dye model trajectories and resulting FRET efficiency vs normalized mean fluorescence lifetime and moments difference for A-B) Isotropic Spring, C-D) Anisotropic Spring, E-F) Elastic Pendulum, as described in Section II. G-H) Sample trajectory of a spherical Brownian motion on the unit sphere $S^{2}$. Such processes are used to model the diffusion of the electric dipole moment. Each of the plots B,D,F,H shows the FRET efficiency vs normalized mean fluorescence lifetime, and the moments difference when the dipole orientation for donor and acceptor is included in the FRET process. Each color in A,C,E,G is an individual burst trajectory. Each case simulates 25000 trajectories over 7 hours. The values in plots B,D,F,H correspond to each simulated state.
In the anisotropic case, illustrated in Figure 4C, we use a diagonal spring matrix with two entries being $pk_{eff}$ and the third being $k_{eff}$ with $p\in[0,1]$. Therefore, the stationary distribution is an ellipse with major axes determined by the entries of the spring matrix. In this section rotational dynamics have not yet been considered; it will be covered in Section II.3.
The two-dimensional dynamics in the anisotropic case can be used to investigate the influence of the orientation of the stationary distribution on the resulting dynamic shift. Such a scenario is exemplified in the case when the planes formed by the major axes of each stationary ellipse are mutually orthogonal. Since the stationary distribution for the isotropic case is a sphere and is perfectly symmetric, this can only arise in the anisotropic case.
Furthermore, these models have the added benefit of having an analytical expression for the inter-dye displacement, especially in the isotropic case. Since the coordinates will be Gaussian distributed the distance between them is simply Rayleigh distributed. [70] This distribution is unimodal, and therefore the only mixing present is due to the variance of the stationary distributions. This mixing is therefore strongly dependent on the flexibility of the dyes.
@@ II.2 Elastic Pendulum Model
Similar to Equation (12), $B_{t}$ is Brownian motion, $\sigma>0$ is the volatility of the random fluctuations that are modeled as Brownian motions, and $k$ is the spring constant. Note that each superscript/subscript is indexed by each of the three components ($r,\theta,\phi$) explained in the next sentences. Importantly, the system is considered in spherical coordinates. The radial dynamics, $r_{t}$ evolve according to the spring dynamics explained in Section II.1, with slight alterations due to the change of coordinates. The angular parts of the motion are given by the standard nonlinear pendulum force in the polar direction $\theta_{t}$ and free diffusion in the azimuthal direction $\phi_{t}$. Figure 4E shows a sample dye trajectory.
@@
Moreover, this model presents a possible explanation for the dynamic shift induced by dye motion due to the non-Gaussian inter-dye distributions, as shown in Figure 5.
@@ II.3 Orientational Dynamics
The final consideration involves the orientational dynamics of the electric dipole moments of the dyes. As discussed in Section I, the Förster radius is dependent on the $\kappa$ parameter, which is dependent on the mutual orientations of the electric dipole moments $\hat{\mu}_{A},\hat{\mu}_{D}$ and the inter-dye displacement unit vector $\mathbf{R}$. Typically, $\kappa^{2}$ is taken as the mean value of $2/3$ when the unit vectors are considered uniformly distributed on the sphere $S^{2}$.[57, 78, 56] This assumption ignores the temporal aspect of the fluorescent process. Since dyes reorient on timescales faster than fluorescent lifetimes, the energy exchange rate changes during the FRET process. This changes the original CTMC model to a time-inhomogeneous CTMC, and thus, the transfer rates are dependent on the time integral of the infinitesimal transfer rates. [63, 64, 65, 62]
@@
$\begin{cases}\mathbf{I}d\omega_{t}+\omega_{t}\times\mathbf{I}\omega_{t}=-\nu\omega_{t}+d\mathbf{W_{t}}\\ \omega_{t}=d\mathbf{\Phi_{t}}\end{cases}$ (14)
where $\omega$ is angular velocity, $\nu$ is the dynamic viscosity of the surrounding fluid, $dW_{t}$ is a spherical Brownian motion and $\mathbf{\Phi_{t}}$ is the angular position vector. Assuming the dye is overdamped and hence $d\omega_{t}=0$, one obtains the simplified equations
$\begin{cases}\omega_{t}\times\mathbf{I}\omega_{t}=-\nu\omega_{t}+d\mathbf{W_{t}}\\ \omega_{t}=d\mathbf{\Phi_{t}}.\end{cases}$ (15)
@@
$\begin{cases}d\theta_{t}&=\frac{\sigma_{\theta}^{2}}{\tan{(\theta_{t})}}dt+\sigma_{\theta}\circ dB_{t}\\ d\mathbf{\Phi_{t}}&=\frac{\sigma_{\phi}}{\sin{\theta_{t}}}\circ dB_{t}\end{cases}$ (17)
The rotational diffusion coefficients depend on the hydrodynamic radius of the dye $R_{h}$ by the classical relation $D=kT/(8\pi\nu R_{h}^{3})$, where $kT$ denotes the product of the Boltzmann constant and the temperature. A sample trajectory is shown in Figure 4G.
Note that the stationary distribution for such a system is the uniform distribution, providing an ideal starting stochastic process to test the time-dependent behavior of orientational dynamics.[81, 85] The key idea is that the excitation of the fluorophores provides a single initial $\kappa$ value. The relaxation effects are the object of interest, especially with regard to lifetime duration. The notion that $\kappa$ may be close to 00 during the entire FRET process for one excitation but higher for another in the same sampling time provides an additional source of variance in the lifetime distribution. Every FRET process can lead to a different equilibrium, which should depend on the rotational diffusion of the dipole moment, with faster reorientation causing an averaging out effect as mentioned in [57].